OpenStudy (anonymous):

3. 7(5 * 6) = 7 + 5 + 6 18 (7 * 5) * 6 4. m * 1 = o m m + 1 5. 0 – 897 = -897 1 -0

OpenStudy (lifeisadangerousgame):

So, is this how the problems are? 3. 7(5 * 6) a) 7 + 5 + 6 b) 18 c) (7 * 5) * 6 4. m * 1 = a) o b) m c) m + 1 5. 0 - 897 = a) - 897 b) 1 c) 0 Right?

OpenStudy (anonymous):

yes. i did them myself already and i ave the answers to them i just wanna check if im correct. i can post them if you want

OpenStudy (lifeisadangerousgame):

Okay sure, let me see your answers

OpenStudy (anonymous):

3) C 4)A 5) B

OpenStudy (anonymous):

@LifeIsADangerousGame

OpenStudy (lifeisadangerousgame):

4 and 5 are incorrect. what's 4 * 1? It's 4. Sometimes variables throw us for a loop. Anything times 1 is still the same number. 5 * 1 is 5, 6 * 1 is 6, and m * 1 is still m. :) 5. 0 - 897 is actually -897. You're subtracting a number from 0, which is nothing, so it becomes a negative number.